# HOPE: A Reinforcement Learning-based Hybrid Policy Path Planner for Diverse Parking Scenarios

### v2

@@ Appendix A Implementation of information fusion transformer
We employ a multi-layer perceptron model for vector-based input to encode it into a feature token. For image-based input, we use a convolution neural network (CNN) with residual blocks to extract 2D features, which are later flattened to match the dimensions of other tokens. Tokens from different views are then fused in the transformer encoder, and their values are selected as outputs. Finally, an MLP-based network is used to map the concatenation of the values to the target output, namely the action for actor-net or the estimated value for value-net. Moreover, we use an auto-encoder structure to train the encoder in self-supervision. The CNN encoder extracts the features from the original image, and a deconvolutional neural network is used to reconstruct the image from the features. The pre-training data is collected with a naive agent trained initially without image input. The loss function of the auto-encoder using a mean square error can be expressed as:
@@
Appendix B Experiments on the Extended Reeds-Shepp Curves
Appendix C Reward function
@@
C-1 Intersection-of-union(IOU) reward
@@ C-2 Distance reward
$D(t)$ is the distance from the vehicle to the target position at time step $t$. $D_{min}$ is a hyperparameter to avoid too large reward value.
C-3 Time consumption penalty
@@
C-4 Overall reward function
The above basic rewards and step rewards are linearly combined to form the final reward function:
@@ Appendix D Hyperparameters
Appendix E Real-world experiment setups
For sensors, we used Hesai Pandar40, RoboSense RS-LiDAR-32 and Bpearl LiDAR to support the vehicle’s localization and perception modules and no GPS was included. Figure 8 illustrates the framework of the parking system and how our method operates in real-world scenarios. The input to the planning module, where our method is applied, consists of a BEV grid map, obstacle distance, and the position of the parking spot. The target parking spot is selected by the driver and its precise location is obtained from a high-definition map, then transformed into the ego coordinate system through the localization system. Obstacle distance information is derived from the LiDAR point cloud, while the BEV grid map is generated by augmenting the drivable area from the high-definition map with real-time detected obstacles from the LiDAR. To mitigate the impact of upstream perception errors, we also add virtual occupancy grids in other parking spaces in the BEV grid map based on high-definition map information, ensuring the vehicle does not drive into these areas during maneuvering. These inputs are then converted into the format required by the network, which outputs a one-step planned path. Finally, a Quadratic Programming (QP) and Model Predictive Control (MPC) approach is used for path following to obtain the vehicle control commands from the waypoints. Besides, we also utilize a collision avoidance module to give way to dynamic obstacles in practice. All the perception and localization modules along with our planning algorithm are deployed on an Intel i7-1165G7 2.80GHz CPU.
@@ Appendix E Real-world experiment setups
Appendix F Proof of Convergence for the Hybrid Policy

### v3

@@ Appendix A Implementation detials
For the implementation of the information fusion transformer, we employ a multi-layer perceptron model for vector-based input to encode it into a feature token. For image-based input, we use a convolution neural network (CNN) with residual blocks to extract 2D features, which are later flattened to match the dimensions of other tokens. Tokens from different views are then fused in the transformer encoder, and their values are selected as outputs. Finally, an MLP-based network is used to map the concatenation of the values to the target output, namely the action for actor-net or the estimated value for value-net. Moreover, we use an auto-encoder structure to train the encoder in self-supervision. The CNN encoder extracts the features from the original image, and a deconvolutional neural network is used to reconstruct the image from the features. The pre-training data is collected with a naive agent trained initially without image input. The loss function of the auto-encoder using a mean square error can be expressed as:
@@
Appendix B Reward function
@@
B-1 Intersection-of-union(IOU) reward
@@ B-2 Distance reward
$D(t)$ is the distance from the vehicle to the target position at time step $t$. $D_{min}$ is a hyperparameter to avoid large rewards.
B-3 Time consumption penalty
@@
B-4 Overall reward function
The above basic rewards and step rewards are linearly combined to form the final reward:
@@ B-4 Overall reward function
Appendix C Real-world experiment setups
For sensors, we used Hesai Pandar40, RoboSense RS-LiDAR-32 and Bpearl LiDAR to support the vehicle’s localization and perception modules and no GPS was included. Figure 7 illustrates the framework of the parking system and how our method operates in real-world scenarios. The input to the planning module, where our method is applied, consists of a BEV grid map, obstacle distance, and the position of the parking spot. The target parking spot is selected by the driver and its precise location is obtained from a high-definition map, then transformed into the ego coordinate system through the localization system. Obstacle distance information is derived from the LiDAR point cloud, while the BEV grid map is generated by augmenting the drivable area from the high-definition map with real-time detected obstacles from the LiDAR. To mitigate the impact of upstream perception errors, we also add virtual occupancy grids in other parking spaces in the BEV grid map based on high-definition map information, ensuring the vehicle does not drive into these areas during maneuvering. These inputs are then converted into the format required by the network, which outputs a one-step planned path. Finally, a Quadratic Programming (QP) and Model Predictive Control (MPC) approach is used for path following to obtain the control commands from the waypoints. Besides, we also utilize a collision avoidance module to give way to dynamic obstacles in practice. The vehicle only follows a path if it has been successfully planned as collision-free and is connected to the target parking position. All modules and our algorithm are deployed on an Intel i7-1165G7 2.80GHz CPU.
@@ Appendix C Real-world experiment setups
Appendix D Experiments on the Extended Reeds-Shepp Curves
Figure 8: The success frequency of different RS curve types
Appendix E Proof of Convergence for the Hybrid Policy